OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. R(3, 3), S(-6, -6)

OpenStudy (anonymous):

Do you know how to get the slope?

OpenStudy (anonymous):

No I dont That is what I need help with

OpenStudy (anonymous):

okay to find the slope you use the equation \[y ^{2}-y ^{1}\div x ^{2}-x ^{1}\]

OpenStudy (anonymous):

can you do that for me?

OpenStudy (anonymous):

yes give me a minute

OpenStudy (anonymous):

-6-3 divided by -6-3

OpenStudy (anonymous):

yes so the slope is 1, now use the formula \[y-y ^{1}=m(x-x ^{1})\]

OpenStudy (anonymous):

-6-3=-9m(-6-3=-9

OpenStudy (anonymous):

actually it is \[y-3=1(x-3) \] solve for slope intercept form, do you know what that is?

OpenStudy (anonymous):

i am sorry i dont i have a really hard time in math but i am willing to figure it out

OpenStudy (anonymous):

okay lets start from the beginning, You are trying to find the standard form equation of a line running in between the points, (3,3) & (-6,-6). Standard form is Ax+By=C and the equation i gave you before is to get y=mx+b, which is the equation you get before standard form. Are you good so far?

OpenStudy (anonymous):

Yes I am

OpenStudy (anonymous):

So you using the equation \[y-3=1(x-3)\] becomes y-3=1x-3. Now isolate y by itself and what do you get?

OpenStudy (anonymous):

Y-3

OpenStudy (anonymous):

Your answer becomes y=1x+0, because you add 3 to both sides to get y on one side.

OpenStudy (anonymous):

so that is the answer?

OpenStudy (anonymous):

Now to get standard form, Ax+By=C. Subtract 1x from each side to get -1x+y=0 (change all the signs when the leading term is negative) Your answer is x-y=0